OpenStudy (destinykiara99):

State the domain of the rational function f(x) = 6 DIVIDED BY 4-x

jimthompson5910 (jim_thompson5910):

this is the function \[\Large f(x) = \frac{6}{4-x}\] right?

OpenStudy (destinykiara99):

yes

jimthompson5910 (jim_thompson5910):

recall that you cannot divide by zero if 4-x were equal to zero, then 4-x = 0 x = ???

jimthompson5910 (jim_thompson5910):

tell me what you get when you solve 4-x = 0 for x

OpenStudy (destinykiara99):

Openstudy keeps kicking me off, sorry. x=4

jimthompson5910 (jim_thompson5910):

yes, so if x = 4, then the denominator is 0

jimthompson5910 (jim_thompson5910):

we must kick x = 4 out of the domain. The value of x can be any other number you want

OpenStudy (destinykiara99):

The answer would be "All real numbers except 4" then, correct?

jimthompson5910 (jim_thompson5910):

`The answer would be "All real numbers except 4" then, correct?` that is correct